OpenStudy (anonymous):

3(x-1)^2 =27 Solve quadratic with square roots

OpenStudy (anonymous):

Hi! So this question is worded funny... but it's really just asking you to solve for x in the easiest way possible.

OpenStudy (anonymous):

I think the answer is x=2

OpenStudy (anonymous):

Nope, but a good guess. The easiest approach would be to solve using algebra. What would your first step be if that is the case?

OpenStudy (anonymous):

Here's a hint.... If you first divided both sides by 3, then would it be possible to take the square root of both sides?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

3(x+1)^2=27 (x+1)^2=9 (x+1)=9^1\2 x+1=3 x=3-1 x=2